OpenStudy (anonymous):

the derivative of P=.6P(t) and P(0)=1.....How large is the population after 2 years? and How quickly is the population growing after 2 years?

OpenStudy (anonymous):

for the first one I put e^1.2.,...but wrong

OpenStudy (unklerhaukus):

\[\lambda P(t)=\frac{\text d P(t)}{\text d t}\]

OpenStudy (anonymous):

yes so this makes the equation e^(.6t) and then wouldnt i just plug in 2 for t to find out how large the population is after 2 years

OpenStudy (unklerhaukus):

sounds right to me

OpenStudy (anonymous):

so when i put in e^1.2 it says its wrong

OpenStudy (anonymous):

or if you dont know what about the second question

OpenStudy (unklerhaukus):

so lambda is 0.6

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[P(t)=P_0 e^{\lambda t}\] \[P(2\text {yr})=P_0 e^{\lambda \cdot 2\text {yr}}\] \[P(2\text {yr})= e^{0.6 \cdot 2\text {yr}}= e^{1.2}\approx3.320=3\quad (\text{natural number })\]

OpenStudy (unklerhaukus):

the population has to be natural number right?